# CLLMs: Consistency Large Language Models

### v3

@@ 3.2.1 Jacobi Trajectory Collection
  Input: prompt set $\mathcal{O}$, n-gram size $n$, max new tokens $N$, target LLM $p$
@@ 3.2.2 Training
  Input: Jacobi trajectory dataset $\mathcal{D}$, n-gram size $n$, the weight factor $\omega$, CLLM $q_{\theta}(\cdot|{\bm{x}})$
@@ 4.3 Ablation Studies
Different lengths of $n$-token sequence. We investigate how different $n$-token sequence lengths in the Jacobi trajectory dataset affect CLLMs’ performance on GSM8K. We employ varying lengths to generate the Jacobi dataset and train the CLLMs accordingly. Figure 3 illustrates that CLLMs consistently maintain generation quality while the models are trained with different lengths. In practice, longer sequence lengths come at cost of increased computational overhead during inference. In Figure 3, significant degradation inference speed can thus be observed when the when the $n$-token sequence length exceeds 64.
@@ 4.4 Limitations and Discussion
In our experiments, we observe that achieving significant speedup while maintaining good generation quality with a CLLM relies strongly on having a high-quality Jacobi trajectory dataset. Therefore, data cleaning is crucial, as discussed in Section 3.2.1. Dataset size also plays a role as described in Section 3 and shown in Table 4, although to a lesser extent. For instance, Jacobi trajectories generated with only 10% of the Code-Search-Net Python dataset is able to yield a 2.9×2.9$\times$2.9 × speedup as demonstrated in Table 2. However, for open-domain datasets like ShareGPT, more data is necessary for improved efficiency.
@@ Appendix B Comparison with Baseline Algorithms
Table 7: All speedups are relative to the vanilla AR. CLLMs has the best memory efficiency and adaptability as it requires no modifications to the model. $*{}^{*}$ refers to capability of achieving more than 3×3$\times$3 × speedup on at least one of our benchmarks. Jacobi decoding doesn’t always lead to a speedup as discussed in Section 3.1, so we denote it with yes.
@@
Eagle yes no no no no $*{}^{*}$
Medusa no no no no no $*{}^{*}$
CLLMs ( Ours ) no no yes yes yes $*{}^{*}$

### v4

@@ 3.2.1 Jacobi Trajectory Collection
  Input: prompt set $\mathcal{O}$, n-token sequence size $n$, max new tokens $N$, target LLM $p$
@@ 3.2.2 Training
  Input: Jacobi trajectory dataset $\mathcal{D}$, n-token sequence size $n$, the weight factor $\omega$, CLLM $q_{\theta}(\cdot|{\bm{x}})$
@@ 4.3 Ablation Studies
In this section, we evaluate the impact of various hyperparameter selections on the performance of CLLMs.
@@
Different lengths of $n$-token sequence. We investigate how different $n$-token sequence lengths in the Jacobi trajectory dataset affect CLLMs’ performance on GSM8K. We employ varying lengths to generate the Jacobi dataset and train the CLLMs accordingly. Figure 3 illustrates that CLLMs consistently maintain generation quality while the models are trained with different lengths. In practice, longer sequence lengths come at cost of increased computational overhead during inference. In Figure 3, significant degradation inference speed can thus be observed when the $n$-token sequence length exceeds 64.
@@ 4.4 Limitations and Discussion
In our experiments, we observe that achieving significant speedup while maintaining good generation quality with a CLLM relies strongly on having a high-quality Jacobi trajectory dataset. Therefore, data cleaning is crucial, as discussed in Section 3.2.1. Dataset size also plays a role as described in Section 3 and shown in Table 4, although to a lesser extent. For instance, Jacobi trajectories generated with only 10% of the Code-Search-Net Python dataset is able to yield a 2.9×2.9$\times$2.9 × speedup as demonstrated in Table 2. However, for open-domain datasets like ShareGPT, more data is necessary for improved efficiency. The computation cost for CLLMs training is moderate and discussed in Appendix D.
@@ 5 Conclusion
Acknowledgments
This work was supported by Key R&D Program of Shandong Province, China (2023CXGC010112), NSF of China (Nos. 62306176, 62102257), National Key R&D Program of China (2022YFB4500200), Natural Science Foundation of Shanghai (No. 23ZR1428700), and CCF-Baichuan-Ebtech Foundation Model Fund.
@@ Appendix B Comparison with Baseline Algorithms
Table 7: All speedups are relative to the vanilla AR. CLLMs has the best memory efficiency and adaptability as it requires no modifications to the model. yes refers to capability of achieving more than 3×3$\times$3 × speedup on at least one of our benchmarks. Jacobi decoding doesn’t always lead to a speedup as discussed in Section 3.1, so we denote it with yes.
@@
Eagle yes no no no no yes
Medusa no no no no no yes
CLLMs ( Ours ) no no yes yes yes yes
@@ Appendix C Pesudo Code for Jacobi Decoding with KV Cache
Appendix D Computation cost for CLLMs training
For the computation required for dataset generation, the cost is low and it’s a one-time overhead. In the cases where the dataset size is large, for example for CodeSearchNet-Python, only $10\%$ of the dataset is required to generate Jacobi trajectories and the trained CLLMs obtain around 2.5× speedup on average. More details are shown in the table below.
Table 8: Computation required for dataset generation. The estimated generation time is based on sequential generation with batch size = 1. We can further reduce the generation time with serving systems like vLLM (Kwon et al., 2023) with batch size = 16 or more. We give an example of estimated training time with vLLM using batch size = 16 in the table as well. All time is estimated by a single A100 40G GPU*hours.
Dataset $\sharp$ Generated tokens Estimated generation time Estimated generation time(vLLM)
Spider $2\mathrm{M}$ 5 $<1$
GSM8K $10\mathrm{M}$ 14 $\sim 1$
CodeSearchNet-Python $100\mathrm{M}$ 100 8
ShareGPT $200\mathrm{M}$ 120 10
[…]
Table 9: Computation required for consistency training.
Dataset Training time % of pre-training cost Training resources
Spider 2 hours $<0.01\%$ 8 A100 40GB GPUs
GSM8K 12 hours $\sim 0.01\%$ 8 A100 40GB GPUs
CodeSearchNet-Python 22 hours $\sim 0.1\%$ 8 A100 40GB GPUs
ShareGPT 30 hours $\sim 0.2$ 8 A100 40GB GPUs